OpenStudy (anonymous):

The price of a share of Arbor's stock increased from 24 1/2 to 36 1/2. How many points did the stock increase? -12 12 -61 61

OpenStudy (phi):

I would subtract 36 1/2 - 24 1/2 can you do that ?

OpenStudy (anonymous):

12

OpenStudy (phi):

yes

OpenStudy (anonymous):

The price of a share of Alumina stock decreased from 29 to 17 1/4. Which of the following equations represents this statement? 29 + X = 17 1/4 29 - X = 17 1/4 29 X = 17 1/4 X / 29 = 17 1/4

OpenStudy (phi):

this is a bit of a goofy question. but we can guess what they want: decrease means go down, so we should think subtract

OpenStudy (anonymous):

AM LOST HELP ME

OpenStudy (phi):

they are thinking: 29 minus an unknown amount equals 17 1/4 which choice means that ?

OpenStudy (anonymous):

B

OpenStudy (phi):

yes

OpenStudy (anonymous):

A store decreased its price on a computer by $112 to $478. What was the original price? $366 $599 $590 $378